There are 30 people that donated to a church. The amount each person donated has probability density function

,(1)-(*(50-r), (50-x), ifo ifo < x < 50 otherwise TA

Find out the probability that exactly 5 people donated between 20 and 30.
Answer #1

here probability that a random person donate between 20 and 30:

P(20<X<30)=\int_{20}^{30} f(x) dx= \int_{20}^{30} (1/1250)*(50-x) dx =(50x-x2/2)/1250 |3020 =0.20

hence from binomial distribution ;

probability that exactly 5 people donated between 20 and 30=P(X=5)=20C5(0.2)5(0.8)15=0.1746